a) The root mean square displacement for diffusion in the x-direction is <x2> = 201 Use this result to show the rms displacement for diffusion in three dimensions is 6Dt. Hint: Use the Pythagoras equation.
Answer #1

The mean square displacement for diffusion in 3 dimensions is

<r^2>=<x^2+y^2+z^2>=<x^2>+<y^2>+<z^2>

This is the result of the following property of mean

<A+B>=<A>+<B>

Given

\sqrt{<x^2>}=\sqrt{2Dt}

Therefore

<x^2>=2Dt

Similarly, for the other 2 individual dimensions, the result will be

<y^2>=<z^2>=2Dt

Therefore RMS displacement for diffusion in 3 dimensions is

\sqrt{<r^2>}=\sqrt{<x^2>+<y^2>+<z^2>}=\sqrt{2Dt+2Dt+2Dt}=\sqrt{6Dt}